Determine the area under a normal distribution curve with µ = 55 and σ = 7 to the right of x = 68. Be sure to draw a normal curve with the area corresponding to the probability.
Answer #1

7 7 ar88 solution Here u=55, 6=7 68-55 a-4 = for a=68, 2= = 1.86 P[a=68)= P(2>186) =1-P(22186) =1-8-9686=50-0314 55 68 Deres